OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by the equation below. y = x2 - 9x + 18

OpenStudy (anonymous):

x intercpts are found by setting y = 0 and solve. basically factor the right side can you do that?

OpenStudy (anonymous):

Factor what ?

OpenStudy (chris):

this is like the last one, factor the right side of the equation so that when both terms are multiplied together they get the same thing.

OpenStudy (chris):

http://www.purplemath.com/modules/solvquad.htm

OpenStudy (chris):

especially look at this and try it: http://www.purplemath.com/modules/factquad.htm

OpenStudy (chris):

x^2 - 9x + 18 = (x )(x )

OpenStudy (chris):

You want two numbers that multiply to get 18, and add to get -9

OpenStudy (anonymous):

Oh Okay .

OpenStudy (anonymous):

Got It , Thannks

OpenStudy (chris):

np! so what do you get though?

OpenStudy (anonymous):

So Would Get (0, 3) and (0, 6) ?

OpenStudy (chris):

close!

OpenStudy (anonymous):

Where Did I Go Wrong ?

OpenStudy (chris):

x^2 - 9x + 18 = (x - 6 )(x - 3 ), and the rest you did right! except - you mixed up your coordinates when you translated into the answer =)

OpenStudy (chris):

(3,0) and (6,0), not the reverse :p

OpenStudy (chris):

but great work, you did the math part right =)

OpenStudy (anonymous):

Lol Okay !